OpenStudy (anonymous):

What fraction is equal to 48%? A. 12/25 B. 120/25 C. 6/25 D. 24/25

jimthompson5910 (jim_thompson5910):

Rule: x% = x/100 so for example, 37% = 37/100

jimthompson5910 (jim_thompson5910):

once you get it into a fraction using that rule, make sure to reduce it as much as possible

OpenStudy (anonymous):

So the answer would be b

jimthompson5910 (jim_thompson5910):

why B?

jimthompson5910 (jim_thompson5910):

how did you get that?

jimthompson5910 (jim_thompson5910):

another example 36% = 36/100 reduce 36/100 to get 9/25 so 36% = 9/25

OpenStudy (anonymous):

Because x=100

jimthompson5910 (jim_thompson5910):

whatever the percentage is, you place over 100

jimthompson5910 (jim_thompson5910):

12% = 12/100 61% = 61/100 92% = 92/100 etc etc